# Physics meets Topology: Physics-informed topological neural networks for learning rigid body dynamics

### v1

@@ 1 Introduction
Modeling dynamic interactions between rigid bodies is an essential challenge across many scientific and engineering disciplines, including biomechanics, robotics, aerospace engineering, and virtual reality. It is not only essential for understanding complex systems and their dynamics, but it also supports the design of advanced engineered systems with counterfactual reasoning – enabling hypotheses to be tested in realistic virtual environments. When all physical parameters of a system are precisely known, traditional general-purpose simulators, such as MuJoCo mujoco , Bullet bullet , ODE ode , or PhysX physx , can be used to generate plausible predictions about system behavior. However, in many real-world scenarios, key aspects of the environment remain unknown – such as current wind conditions, pressure, temperature, and gravity – or are too complex to model explicitly, including surface texture and wetness, material damping, and fatigue. This challenge is particularly relevant for autonomous systems, which must swiftly react to dynamic objects in their surroundings while operating with limited knowledge of the environment – for example, unfamiliar locations, unpredictable weather conditions, or newly encountered obstacles. Consequently, in situations where numerous aspects of a scene are uncertain, learning-based approaches offer a more practical and effective solution, allowing systems to implicitly learn missing parameters from past observations.
Figure 1: Comparison between graph and combinatorial complex representation. (a) Standard graph representation $\mathcal{G}=(\mathcal{V},\mathcal{E})$ used in meshgraphnet ; fignet ; (b) Our proposed Combinatorial Complex representation $\mathcal{X}$. While the graph $\mathcal{G}$ only supports nodes and edges, the combinatorial complex $\mathcal{X}$ enables us to define a hierarchy of higher-order groups of nodes $X^{(r)}$ (e.g. mesh triangles $X^{(2)}$, collision contacts $X^{(3)}$, and objects $X^{(4)}$) and assign distinct features to them. Only four $X^{(2)}$ cells and one $X^{(3)}$ cell are shown on (b) to preserve visibility.
Furthermore, learning directly from data can bridge the often substantial simulation-to-reality gap. Traditional simulation tools are limited in their accuracy due to the inherently complex nature of rigid body interactions. They necessitate explicit computation of forces, including contact and friction forces, which are notoriously difficult to model accurately challenges_stiff_contact . In contrast, data-driven approaches can implicitly capture intricate and subtle dynamics – such as the fine granularity of object surfaces that significantly affects collisions — thereby offering a more comprehensive and realistic model for physical interactions.
@@
Graph Neural Networks (GNNs) gnn have emerged as a powerful framework for modeling complex relational data by effectively capturing and propagating information across interconnected entities. This capability makes them well-suited for representing the intricate and dynamic interactions inherent in 3D environments. Unlike vision-based methods, which often rely on indirect feature extraction, GNNs provide a more accurate and direct representation of 3D interactions by explicitly modeling the underlying relational structure of the data. Within this framework, particle-based methods graph_network_simulator ; physics_consistent_particles ; subequivariant_gnn ; gns_collisions , which represent each object as a collection of dense particles, have proven effective in simulating liquids, soft materials, simple rigid bodies graph_network_simulator , and contact collisions gns_collisions . However, these methods face significant scalability challenges because their computational cost increases quadratically with scene complexity (e.g. number of objects and collisions). Furthermore, their ability to accurately model collisions is constrained by the particle-based representation, which lacks physical consistency; real-world collisions occur between continuous surfaces, not individual particles.
@@
Recently, novel graph-based methods that operate directly on surfaces have been developed meshgraphnet ; fignet . These approaches transform the meshes of rigid bodies into graph structures composed of vertices and edges, which are then processed using message-passing architectures relational_inductive_biases . Despite their innovative design, these methods face fundamental limitations inherent in basic graph structures, which cannot effectively represent hierarchies or groups of vertices, such as entire surfaces or objects. Restricted to pairwise interactions, GNNs typically lack the expressiveness needed to capture emergent group-level dynamics, which are crucial in rigid body systems where nodes, surfaces, and objects are tightly coupled. Additionally, GNNs face inefficient information propagation: messages can only reach immediate neighbors (1-hop) per layer, requiring many layers to reach distant nodes. This, in turn, introduces issues like over-smoothing gnn_oversmoothing , ultimately reducing model performance.
@@
To partially address those limitations, Allen et. al. fignet proposed duplicating shared object properties across each node and introducing virtual points to facilitate long-range information transfer within objects. Nevertheless, the way neural messages propagate across the graph does not accurately reflect the underlying physical interactions. Specifically, although collisions are computed between mesh triangles, the subsequent message passing occurs in a traditional "node-to-node" fashion – propagating from the points of contact to all other nodes within the object. This approach is inconsistent with Newtonian physics governing inelastic collisions, where energy and momentum are transmitted centrally through the center of mass rather than being distributed solely through individual nodes.
@@
Ensuring consistency between real-world phenomena and their representation in deep neural networks is essential for effectively capturing complex interactions. In the context of rigid body dynamics, this requires improved modeling techniques and a refined processing of object shapes. A promising approach involves leveraging higher-order groups of nodes within graph-based data Battiston2021 ; nature_simplicial ; nature_higher_order , which has found promising applications in domains like chemistry schnet ; Fang2022 ; topologynet ; schutt2021equivariant and biology Pineda2023 . This approach, known as Topological Deep Learning topological_deep_learning ; bodnar_2022 ; topological_deep_learning_survey , has demonstrated promising results in structured data environments scn_trajectory_pred ; cw_networks . However, its application to geometric data remains predominantly limited to static scenarios, primarily focusing on classification tasks while overlooking temporal dynamics and multi-entity interactions. Moreover, existing topological neural networks typically do not take advantage of available physical knowledge about the task in their message-passing mechanisms.
@@
Figure 2: Spatiotemporal combinatorial complexes (a) Real-world observations111 The sequence shown in Fig. 1 is taken from experiment n°1 of the MoVi-A dataset. are represented by (b) spatiotemporal combinatorial complexes $\{\mathcal{X}^{t}\}$. The topology and features of each combinatorial complex evolve over time to accurately represent the environment.
@@
In this article, we propose "Higher-Order topological Physics-informed Network" (HOPNet), a novel and versatile framework designed to (1) model complex dynamic object geometries using higher-order spatiotemporal topological groups, and (2) learn rigid body dynamics through physics-informed topological neural architectures operating on these topological representations. HOPNet effectively addresses the challenges of accurately modeling dynamic rigid body meshes with evolving topological structures by grouping relevant nodes and establishing a clear hierarchy among nodes, mesh triangles, and objects. Learnable messages are exchanged between these higher-order groups using pathways specifically defined based on Newtonian principles, thereby integrating physics-based knowledge directly into the learning process. We evaluate our framework on tasks involving forward dynamics and collisions between rigid bodies in increasingly complex scenarios. HOPNet demonstrates high accuracy with complex shapes and multi-object collisions, successfully capturing highly nonlinear and abrupt dynamics while maintaining its accuracy over long rollouts. Additionally, the method exhibits strong generalization to both unseen and counterfactual scenarios, indicating its effectiveness in learning the underlying fundamental physical laws. This counterfactual capability enables rigorous exploration of possible outcomes across various hypothetical conditions, reinforcing our framework’s utility in diverse scientific and engineering domains. Finally, our framework’s learning efficiency and robustness to hyperparameter selection further minimize the need for fine-tuning, supporting broad applicability and ease of deployment across different tasks.
@@ 2.1 Object meshes as spatiotemporal topological groups
Rigid-body simulations represent objects as meshes, where each mesh $M$ is defined by a set of node positions $\{\bm{x}_{i}\}_{i=1..N}$ and a set of faces $\{f_{j}\}_{j=1..F}$. The faces describe the connections between nodes, forming the surface of the mesh. Each triangular face $f_{j}$ is defined by three nodes $\{a,b,c\}$, where $a$, $b$, and $c$ are unique node indices within the face. While all meshes in our experiments utilize triangular faces, the proposed framework can generalize to faces with different numbers of nodes (e.g. tetrahedron, hexahedron, etc.).
@@
Figure 3: Overview of our method. (a) Autoregressive rollout approach222 The sequence shown in Fig. 2 is taken from experiment n°2 of the MoVi-A dataset. ; (b) Physics-informed message-passing strategy. Our sequential message-passing is inspired by Newtonian laws and tailored to process collisions. It first processes ongoing collisions (steps 1, 2, and 3) and then updates individual objects (steps 4 and 5). Finally, the per-node $\hat{\ddot{\bm{x}}}_{i}$ and per-object $\hat{\ddot{\bm{x}}}_{n}$ acceleration are computed (step 6) and the final object poses $(\hat{\bm{x}}^{t+1}_{n},\hat{\bm{q}}^{t+1}_{n})$ are obtained (step 7) with shape matching. The embedding vectors of each cell type $X^{(r)}$ of rank $r$ on a combinatorial complex $\mathcal{X}$ are noted $\bm{h}^{(r)}$. Other mathematical symbols are explained in "Methods" and summarized in Supplementary Information Sec 1.
@@
At any given timestep $t$, the entire environment is represented by a single global mesh $M^{t}$ that may contain one or multiple objects. In previous works meshgraphnet ; fignet , this global mesh $M$ was converted into a standard graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where vertices $\mathcal{V}=\{v_{i}\}$ represent the mesh nodes and edges =$\mathcal{E}=\{e_{v_{s}\rightarrow v_{r}}\}$ are bidirectionally created between the nodes of each face (as illustrated in Fig. 1a). However, this approach discards critical information about the scene: (1) cohesion of each individual object: the integrity of each individual object within the global mesh is not preserved, and (2) the consistency of surface triangles: the uniformity and structural integrity of the mesh surfaces are lost. In this uniform representation, all nodes are treated as identical entities, and objects are only distinguishable by the absence of connections between their respective nodes. Consequently, by losing most of the structural information and retaining only the lowest-level connectivity, the representation is reduced to simple node-to-node connections, thereby oversimplifying the true geometry.
@@
To overcome the inherent limitations of previous approaches, our framework introduces a novel representation that preserves both the cohesion of individual objects and the consistency of mesh surfaces. This is achieved by defining meshes as higher-order spatiotemporal topological groups. We propose to represent the global mesh $M$ as a Combinatorial Complex (CC) – a higher-order domain that allows the definition of groups of nodes (called "cells") and introduces a hierarchical structure between them (called "rank" and noted $r$). Unlike other methods meshgraphnet ; fignet ; gns_collisions , our representation enables a more accurate and complete representation of the global mesh $M$ by allowing us to explicitly define objects and mesh triangles. Our proposed representation $\mathcal{X}$ uses cells of five different types, as illustrated in Fig. 1b: nodes $X^{(0)}$, edges $X^{(1)}$, mesh triangles $X^{(2)}$, collision contacts $X^{(3)}$, and objects $X^{(4)}$. Each cell $X^{(r)}$ is assigned specific features $\bm{h}^{(r)}$ – such as velocity for nodes or mass for objects – unlike traditional graph structures, where only nodes and edges are associated with features. This enriched representation captures both geometric and physical properties, improving the model’s ability to handle complex interactions between rigid bodies.
@@
Finally, we represent the evolution of multi-body environments over time using a sequence of spatiotemporal combinatorial complexes $(\mathcal{X}^{t_{0}},\mathcal{X}^{t_{1}},\mathcal{X}^{t_{2}},...)$, illustrated in Fig. 1. Each object in $\mathcal{X}^{t}$ has a specific position and rotation which evolves over time. Our method estimates the next state of the system $\hat{\mathcal{X}}^{t+1}$ from a history of previous states $\{\mathcal{X}^{t-1},\mathcal{X}^{t}\}$. This enables trajectory rollouts by iteratively feeding the model with its previous predictions, generating a sequence such as $\{\mathcal{X}^{t_{0}},\mathcal{X}^{t_{1}},\hat{\mathcal{X}}^{t_{2}},\hat{\mathcal{X}}^{t_{3}},...,\hat{\mathcal{X}}^{t_{k}}\}$. This autoregressive approach is illustrated in Fig. 3a.
@@ 2.2 Physics-informed learnable messages
By introducing higher-order groups within combinatorial complexes, we establish explicit pathways for information flow between cells. Unlike standard message-passing neural architectures relational_inductive_biases used on graphs – where information is typically exchanged only between individual nodes through directed edges – our framework enables learnable messages to be transmitted between any two cells that share at least one common node. This extension allows for efficient information flow not only between nodes but also across groups of nodes, effectively modeling more complex relationships and interactions. Additionally, it supports long-range information transfer without requiring the multiple layers of message passing typically needed in traditional GNNs. Consequently, this enhances the expressiveness, learning efficiency, and accuracy of the learned dynamics.
@@
The incorporation of higher-order groups provides greater flexibility in exchanging information between rigid bodies and their surfaces. Instead of allowing information to flow indiscriminately among all cells, we embed physics-informed inductive biases to define precise pathways for message passing within the combinatorial complexes. This innovative approach offers several key benefits: (1) Reduced computational costs and model complexity: by restricting message passing to targeted pathways, we minimize unnecessary computations and streamline the model architecture;(2) Enhanced learning efficiency: physically-grounded pathways guide the neural network, accelerating the learning process and improving convergence; and (3) Improved model explainability: aligning message flows with physical principles enhances the interpretability of the model, making it easier to understand and trust its predictions.
Figure 4: Autoregressive rollout performance on benchmark datasets. (a) Translation and (b) orientation errors after a rollout horizon of $T=50$ on all dynamic objects in the scene. Error bars indicate the minimum, mean, and maximum performance across three independent random seeds. Comparisons with FIGNet fignet and MeshGraphNet meshgraphnet are based on the results reported in fignet .
@@
To simulate rigid body dynamics, we incorporate Newtonian laws to orchestrate a sequential flow of learnable messages. To effectively capture the abrupt and highly nonlinear dynamics of collisions, we consider two distinct scenarios: (1) Independent evolution of objects: here, an object’s trajectory is only influenced by external forces. Using the principles of conservation of momentum and energy, we compute its acceleration solely based on its current state; and (2) Colliding objects: i scenario, collisions result in the exchange and loss of energy between colliding objects. This requires an additional initial step to process the collision, transfer energy, and update each object’s state before computing their respective accelerations.
@@
We formalize these fundamental principles into a unified, sequential, physics-informed higher-order message-passing framework for Combinatorial Complexes (CCs), as illustrated in Fig. 3b. This novel approach is specifically designed to capture abrupt non-linear dynamics, thereby enhancing the modeling of rigid-body dynamics. First, the features of each triangular face, $\bm{h}^{(2)}$, are enriched by incorporating the features of their constituent nodes, $\bm{h}^{(0)}$, and the corresponding high-level object features, $\bm{h}^{(4)}$ (step 1). In the event of collisions, changes in energy and momentum are computed using collision contacts $X^{(3)}$ (step 2). These collision contacts $X^{(3)}$ are defined based on inter-object distance: when objects come within a certain radius $d_{c}$ of each other, a collision contact cell is created. The computed changes in energy and momentum are then propagated back to the colliding higher-order objects $\bm{h}^{(4)}$ (steps 3 and 4). Finally, predictions at both object and node levels are generated based on the current internal states of each object, $\bm{h}^{(4)}$, and its nodes $\bm{h}^{(0)}$ (steps 5 and 6).
@@
Figure 5: Rollout example on the MoVi-A dataset. (a) Ground truth333 The sequences shown in Fig. 3 are taken from experiment n°6 of the MoVi-A dataset. ; (b) Model predictions. Our model makes very accurate and realistic predictions on the trajectories of rigid bodies despite complex multi-objects collisions and long rollout durations.
HOPNet is designed to predict the forward dynamics of all dynamic objects within a scene. Specifically, HOPNet outputs per-node and per-object accelerations $\hat{\ddot{\bm{x}}}_{i}$, which are then integrated using a second-order forward-Euler method with a timestep $\Delta t=1$ to compute the next positions $\hat{\bm{x}}_{i}^{t+1}$ using the previous and current positions $\bm{x}_{i}^{t-1}$ and $\bm{x}_{i}^{t}$. This integration step is formalized in Eq. 1:
$\hat{\bm{x}}_{i}^{t+1}=\hat{\ddot{\bm{x}}}^{t}_{i}+2\bm{x}_{i}^{t}-\bm{x}_{i}^{t-1}$ (1)

### v3

# Integrating Physics and Topology in Neural Networks for Learning Rigid Body Dynamics

@@ 1 Introduction
Modeling dynamic interactions between rigid bodies is an essential challenge across many scientific and engineering disciplines, including biomechanics, robotics, aerospace engineering, and virtual reality. It is not only essential for understanding complex systems and their dynamics, but it also supports the design of advanced engineered systems with counterfactual reasoning—enabling hypotheses to be tested in realistic virtual environments. When all physical parameters of a system are precisely known, traditional general-purpose simulators, such as MuJoCo mujoco , Bullet bullet , ODE ode , or PhysX physx , can be used to generate plausible predictions about system behavior. However, in many real-world scenarios, key aspects of the environment remain unknown—such as current wind conditions, pressure, temperature, and gravity—or are too complex to model explicitly, including surface texture and wetness, material damping, and fatigue. This challenge is particularly relevant for autonomous systems, which must swiftly react to dynamic objects in their surroundings while operating with limited knowledge of the environment—for example, unfamiliar locations, unpredictable weather conditions, or newly encountered obstacles. Consequently, in situations where numerous aspects of a scene are uncertain, learning-based approaches offer a more practical and effective solution, allowing systems to implicitly learn missing parameters from past observations.
Figure 1: Comparison between graph and combinatorial complex representation. (a) Standard graph representation $\mathcal{G}=(\mathcal{V},\mathcal{E})$ used in meshgraphnet ; fignet ; (b) Our proposed Combinatorial Complex representation $\mathcal{X}$. While the graph $\mathcal{G}$ only supports nodes and edges, the combinatorial complex $\mathcal{X}$ enables us to define a hierarchy of higher-order groups of nodes $X^{(r)}$ (e.g., mesh triangles $X^{(2)}$, collision contacts $X^{(3)}$, and objects $X^{(4)}$) and assign distinct features to them. Only four $X^{(2)}$ cells and one $X^{(3)}$ cell are shown in (b) to preserve visibility.
Furthermore, learning directly from data can bridge the often substantial simulation-to-reality gap. Traditional simulation tools are limited in their accuracy due to the inherently complex nature of rigid body interactions. They necessitate explicit computation of forces, including contact and friction forces, which are notoriously difficult to model accurately challenges_stiff_contact . In contrast, data-driven approaches can implicitly capture intricate and subtle dynamics—such as the fine granularity of object surfaces that significantly affects collisions—thereby offering a more comprehensive and realistic model for physical interactions.
@@
Graph Neural Networks (GNNs) gnn have emerged as a powerful framework for modeling complex relational data by effectively capturing and propagating information across interconnected entities. This capability makes them well-suited for representing the intricate and dynamic interactions inherent in 3D environments. Unlike vision-based methods that rely on indirect feature extraction, GNNs provide a more accurate and direct representation of 3D interactions by explicitly modeling the underlying relational structure of the data. Within this framework, particle-based methods mrowca2018flexible ; propnet ; dpinet ; graph_network_simulator ; physics_consistent_particles ; subequivariant_gnn ; gns_collisions , which represent each object as a collection of dense particles, have proven effective in simulating liquids, soft materials, simple rigid bodies graph_network_simulator , and contact collisions gns_collisions . However, these methods face significant scalability challenges because their computational cost increases quadratically with scene complexity (e.g., number of objects and collisions). Furthermore, their ability to accurately model collisions is constrained by the particle-based representation, which lacks physical consistency; real-world collisions occur between continuous surfaces, not individual particles.
@@
Recently, graph-based methods that operate directly on surfaces have been developed meshgraphnet ; fignet . These approaches transform the meshes of rigid bodies into graph structures composed of vertices and edges, which are then processed using message-passing architectures relational_inductive_biases . Several of these methods introduce surface- or object-level representations mrowca2018flexible ; meshgraphnet ; fignet using heuristics such as virtual nodes fignet or hierarchical clustering mrowca2018flexible . While such strategies improve upon particle-based models, they remain limited in how explicitly they encode the underlying physical and geometric structure. Although these methods introduce higher-level concepts, their internal computations still primarily rely on node-level message passing, which may limit their ability to capture emergent group-level dynamics—crucial in rigid body systems where nodes, surfaces, and objects are tightly coupled. Additionally, GNNs face inefficient information propagation: messages can only reach immediate neighbors (1-hop) per layer, requiring many layers to reach distant nodes. This, in turn, introduces issues like over-smoothing gnn_oversmoothing , ultimately making training more challenging and reducing model performance.
@@
To partially address those limitations, Allen et. al. fignet proposed duplicating shared object properties across each node and introducing virtual points to facilitate long-range information transfer within objects. Nevertheless, the way neural messages propagate across the graph does not accurately reflect the underlying physical interactions. Specifically, although collisions are computed between mesh triangles, the subsequent message passing occurs in a traditional "node-to-node" fashion—propagating from the points of contact to all other nodes within the object. This approach is inconsistent with Newtonian physics governing inelastic collisions, where energy and momentum are transmitted centrally through the center of mass rather than being distributed solely through individual nodes.
@@
Ensuring consistency between real-world phenomena and their representation in deep neural networks is essential for effectively capturing complex interactions. In the context of rigid body dynamics, this requires improved modeling of rigid bodies and a refined processing of collisions. A promising direction involves leveraging higher-order groups of nodes within graph-based data Battiston2021 ; nature_simplicial ; nature_higher_order , which has found promising applications in domains like chemistry schnet ; Fang2022 ; topologynet ; schutt2021equivariant and biology Pineda2023 . This approach, known as Topological Deep Learning topological_deep_learning ; bodnar_2022 ; topological_deep_learning_survey , has demonstrated promising results in structured data environments scn_trajectory_pred ; cw_networks . However, its application to geometric data remains predominantly limited to static scenarios, primarily focusing on classification tasks while overlooking temporal dynamics and multi-entity interactions. Moreover, existing topological neural networks typically do not incorporate physical knowledge into their message-passing mechanisms.
@@
Figure 2: Spatiotemporal combinatorial complexes (a) Real-world observations are represented by (b) spatiotemporal combinatorial complexes $\{\mathcal{X}^{t}\}$. The topology and features of each combinatorial complex evolve over time to accurately represent the environment. The illustrated sequence is taken from experiment 1 of the MOVi-A dataset.
@@
In this article, we propose "Higher-Order topological Physics-informed Network" (HOPNet), a novel and versatile learning-based message-passing framework designed to (1) model complex dynamic object geometries using higher-order spatiotemporal topological groups, and (2) learn rigid body dynamics through physics-informed topological neural architectures operating on these topological representations. Unlike prior approaches that heuristically approximate object structure, HOPNet formally encodes surfaces and objects using evolving topological structures that group relevant nodes and establish a clear hierarchy among nodes, mesh triangles, and objects. Learnable messages are exchanged between these higher-order groups along pathways specifically defined based on Newtonian principles, thereby integrating physics-based knowledge directly into the learning process. We evaluate our framework on tasks involving forward dynamics and collisions between rigid bodies in increasingly complex scenarios. HOPNet demonstrates high accuracy with complex shapes and multi-object collisions, successfully capturing highly nonlinear and abrupt dynamics while maintaining its performance over long rollouts. Additionally, the method exhibits strong generalization to both real-world geometries and counterfactual scenarios, indicating its effectiveness in learning the underlying fundamental physical laws. This counterfactual capability enables rigorous exploration of possible outcomes under various hypothetical conditions, reinforcing our framework’s utility in diverse scientific and engineering domains. Finally, our framework’s learning efficiency and robustness to hyperparameter selection further minimize the need for fine-tuning, supporting broad applicability and ease of deployment across different tasks.
@@ 2.1 Object meshes as spatiotemporal topological groups
Rigid-body simulations represent objects as meshes, where at a given timestep $t$, each mesh $M^{t}$ is defined by a set of node positions $\{\mathbf{x}_{i}\}_{i=1..N}$ and a set of faces $\{f_{j}\}_{j=1..F}$. The faces describe the connections between nodes, forming the surface of the mesh. Each triangular face $f_{j}$ is defined by three nodes $\{a,b,c\}$, where $a$, $b$, and $c$ are unique node indices within the face. While all meshes in our experiments utilize triangular faces, the proposed framework can generalize to faces with different numbers of nodes (e.g., tetrahedron, hexahedron, etc.).
@@
Figure 3: Overview of our method. (a) Autoregressive rollout approach; (b) Physics-informed message-passing strategy. Our sequential message-passing is inspired by Newtonian laws and tailored to process collisions. It first processes ongoing collisions (steps 1, 2, and 3) and then updates individual objects (steps 4 and 5). Finally, the per-node $\hat{\ddot{\mathbf{x}}}_{i}$ and per-object $\hat{\ddot{\mathbf{x}}}_{k}$ accelerations are computed (step 6) and the final object poses $(\text{positions }\hat{\mathbf{x}}^{t+1}_{k},\text{quaternionic orientations }\hat{\mathbf{q}}^{t+1}_{k})$ are obtained (step 7) with shape matching. The embedding vectors of each cell type $X^{(r)}$ of rank $r$ on a combinatorial complex $\mathcal{X}$ are noted $\mathbf{h}^{(r)}$. Other mathematical symbols are explained in "Methods" and summarized in Supplementary Information Section 1. The illustrated sequence is taken from experiment 2 of the MOVi-A dataset.
@@
At any given timestep $t$, the entire environment is represented by a single global mesh $M^{t}$ that may contain one or multiple objects. In previous works meshgraphnet ; fignet , this global mesh $M^{t}$ was converted into a standard graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where vertices $\mathcal{V}=\{v_{i}\}$ represent the mesh nodes and edges =$\mathcal{E}=\{e_{v_{s}\rightarrow v_{r}}\}$ are bidirectionally created between the nodes of each face (as illustrated in Figure 1a). However, this approach discards critical information about the scene: (1) cohesion of each individual object: the integrity of each individual object within the global mesh is not preserved, and (2) the consistency of surface triangles: the uniformity and structural integrity of the mesh surfaces are lost. In this uniform representation, all nodes are treated as identical entities, and objects are only distinguishable by the absence of connections between their respective nodes. Consequently, by losing most of the structural information and retaining only the lowest-level connectivity, the representation is reduced to simple node-to-node connections, thereby oversimplifying the true geometry.
@@
To overcome the inherent limitations of previous approaches, our framework introduces a novel representation that preserves both the cohesion of individual objects and the consistency of mesh surfaces. This is achieved by defining meshes as higher-order spatiotemporal topological groups. We propose to represent the global mesh $M^{t}$ as a Combinatorial Complex (CC)—a higher-order domain that allows the definition of groups of nodes (called "cells") and introduces a hierarchical structure between them (called "rank" and denoted $r$). Unlike other methods meshgraphnet ; fignet ; gns_collisions , our representation allows for an explicit, physics-consistent encoding of both mesh triangles and entire objects of the global mesh $M^{t}$, enabling a more faithful modeling of real-world dynamics. Our proposed representation $\mathcal{X}$ uses cells of five different types, as illustrated in Figure 1b: nodes $X^{(0)}$, edges $X^{(1)}$, mesh triangles $X^{(2)}$, collision contacts $X^{(3)}$, and objects $X^{(4)}$. Each cell $X^{(r)}$ is assigned specific features $\mathbf{h}^{(r)}$—such as velocity for nodes or mass for objects—unlike traditional graph structures, where only nodes and edges are associated with features. This enriched representation captures both geometric and physical properties, improving the model’s ability to handle complex interactions between rigid bodies.
@@
Finally, we represent the evolution of multi-body environments over time using a sequence of spatiotemporal combinatorial complexes $\{\mathcal{X}^{t_{0}},\mathcal{X}^{t_{1}},\mathcal{X}^{t_{2}},...\}$, illustrated in Figure 2. Each object in $\mathcal{X}^{t}$ has a specific position and rotation which evolves over time. Our method estimates the next state of the system $\hat{\mathcal{X}}^{t+1}$ from a history of previous states $\{\mathcal{X}^{t-1},\mathcal{X}^{t}\}$. This enables trajectory rollouts by iteratively feeding the model with its previous predictions, generating a sequence such as $\{\mathcal{X}^{t_{0}},\mathcal{X}^{t_{1}},\hat{\mathcal{X}}^{t_{2}},\hat{\mathcal{X}}^{t_{3}},...,\hat{\mathcal{X}}^{t_{k}}\}$. This autoregressive approach is illustrated in Figure 3a.
Figure 4: Autoregressive rollout performance on benchmark datasets. (a) Translation and (b) orientation errors after a rollout horizon of $T=50$ on all dynamic objects in the scene. Error bars indicate the mean and standard deviation across three independent random seeds. Comparisons with FIGNet fignet and MeshGraphNet meshgraphnet are based on the results reported in fignet .
@@ 2.2 Physics-informed learnable messages
By introducing higher-order groups within combinatorial complexes, we establish physically-motivated pathways for information flow between cells. Unlike standard message-passing neural architectures relational_inductive_biases used on graphs—where information is typically exchanged only between individual nodes through directed edges—our framework allows learnable messages can be transmitted between any two cells that share at least one common node. This extension enables efficient information flow not only between nodes, but also between mesh elements and object-level groups, effectively modeling more complex relationships and interactions. Additionally, it supports long-range information transfer without requiring the multiple layers of message-passing typically needed in traditional GNNs. Consequently, this enhances the expressiveness, learning efficiency, and accuracy of the learned dynamics.
@@
The incorporation of higher-order groups provides greater flexibility in exchanging information between rigid bodies and their surfaces. Rather than allowing information to flow indiscriminately among all cells, we embed physics-informed inductive biases to define precise pathways for message-passing within the combinatorial complexes. This innovative approach offers several key benefits: (1) Reduced computational costs and model complexity: by restricting message-passing to targeted pathways, we minimize unnecessary computations and streamline the model architecture;(2) Enhanced learning efficiency: physically-grounded pathways guide the neural network, accelerating the learning process and improving convergence; and (3) Improved model explainability: aligning message flows with physical principles enhances the interpretability of the model, making it easier to understand and trust its predictions.
@@
To simulate rigid body dynamics, we incorporate Newtonian laws to orchestrate a sequential flow of learnable messages. These messages are computed using MultiLayer Perceptrons (MLPs) with learnable weights, which are optimized via gradient descent to capture the underlying physics of motion and collisions. To effectively capture the abrupt and highly nonlinear dynamics of collisions, we consider two distinct scenarios: (1) Independent evolution of objects: here, an object’s trajectory is only influenced by external forces. Using the principles of conservation of momentum and energy, we compute its acceleration solely based on its current state; and (2) Colliding objects: in this scenario, collisions result in the exchange and loss of energy between colliding objects. This requires an additional initial step to process the collision, transfer energy, and update each object’s state before computing their respective accelerations.
@@
We formalize these fundamental principles into a unified, sequential, physics-informed higher-order message-passing framework for Combinatorial Complexes (CCs), as illustrated in Figure 3b. This novel approach is specifically designed to capture abrupt nonlinear dynamics, thereby enhancing the modeling of rigid-body dynamics. First, the features of each triangular face, $\mathbf{h}^{(2)}$, are enriched by incorporating the features of their constituent nodes, $\mathbf{h}^{(0)}$, and the corresponding high-level object features, $\mathbf{h}^{(4)}$ (step 1). In the event of collisions, changes in energy and momentum are computed using collision contacts $X^{(3)}$ (step 2). These collision contacts $X^{(3)}$ are defined based on inter-object distance: when objects come within a certain radius $d_{c}$ of each other, a collision contact cell is created. This mechanism follows a standard practice in neural physics simulators dpinet ; propnet ; fignet , where connectivity is dynamically updated based on spatial proximity. The computed changes in energy and momentum are then propagated back to the colliding higher-order objects $\mathbf{h}^{(4)}$ (steps 3 and 4). Finally, predictions at both object and node levels are generated based on the current internal states of each object, $\mathbf{h}^{(4)}$, and its nodes $\mathbf{h}^{(0)}$ (steps 5 and 6).
@@
Figure 5: Rollout example on the MOVi-A dataset. (a) Ground truth11footnotemark: 1 ; (b) Model predictions. Our model makes very accurate and realistic predictions on the trajectories of rigid bodies despite complex multi-objects collisions and long rollout durations. The illustrated sequence is taken from experiment 6 of the MOVi-A dataset.
HOPNet is designed to predict the forward dynamics of all dynamic objects within a scene. During training, HOPNet outputs per-node $\hat{\ddot{\mathbf{x}}}_{i}$ and per-object accelerations $\hat{\ddot{\mathbf{x}}}_{k}$, which are compared to ground truth values to optimize the model weights.
During inference, only the per-node accelerations are integrated to compute future node positions using a second-order forward-Euler method with a timestep $\Delta t=1$. Given the previous and current positions $\mathbf{x}_{i}^{t-1}$ and $\mathbf{x}_{i}^{t}$, the estimated next node positions $\hat{\mathbf{x}}_{i}^{t+1}$ are obtained as follows in Equation 1:
$\hat{\mathbf{x}}_{i}^{t+1}=\hat{\ddot{\mathbf{x}}}^{t}_{i}+2\mathbf{x}_{i}^{t}-\mathbf{x}_{i}^{t-1}$ (1)
Finally, the object-level positions $\{\hat{\mathbf{x}}^{t+1}_{k}\}_{k=1...K}$ and quaternionic orientations $\{\hat{\mathbf{q}}^{t+1}_{k}\}_{k=1...K}$ are determined using shape-matching shape_matching , which deterministically computes the optimal object positions and orientations fitting the updated node positions, ensuring consistent rigid body motion. A quaternion $q=(w,x,y,z)$ efficiently represents an object’s 3D orientation without ambiguity, unlike Euler angles. More details are available in the "Methods" section.